5x+15y=35
2x+8y=18 using substitution method

Answer by rajarn_shahi456(1)

5x+15y=35
2x+8y=18 using substitution method
the solution of this equation is given below
5x+15y=35
5x=35-15y
x=(35-15y)/5........(i)
Now putting the value of equation (i) in equation 2nd
2((35-15y)/5))+8y=18
((70-30y)/5))+8y=18
(70-30y+40y)/5=18
70+5y=90
5y=20
y=4.
Now puttiing the value of y in equation (i) we get
x=(35-15y)/5
x=(35-15*4)/5
x=-25/5
x=-5
there fore x=-5 and y=4 ans.....